OpenStudy (kikilandry):

how do you convert 100ft per second to miles per hour

OpenStudy (johnweldon1993):

There are 5280 ft in a mile and there are 3600 seconds in an hour So we have \[\large 100\frac{ft}{s} \times (\frac{1 mile}{5280 ft})\times (\frac{3600s}{1hour})=?\]

OpenStudy (kikilandry):

68

563blackghost (563blackghost):

Got the same as well :)

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) \(\Large{By:thomaster}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct

OpenStudy (jonathan34):

need help? still?

OpenStudy (mathmale):

Johnweldon's \[\large 100\frac{ft}{s} \times (\frac{1 mile}{5280 ft})\times (\frac{3600s}{1hour})\] is well done (congrats, johnwelcon). He is using 2 diffrerent conversion factors here, one to convert feet to miles and another to convert seconds to hours. Complete this work by cancelling units of measurement (not all of them!).